OpenStudy (anonymous):

factor the trinomial by grouping: 12m^2-11m+5 help!

OpenStudy (mathmate):

The given expression has complex factors, and grouping is not intuitive. Does the question happen to be \(12m^2-17m+5\)? If it is, then you can group as follows: \(12m^2-12m -5m+5\)

OpenStudy (anonymous):

no. thats why im confused also!

OpenStudy (anonymous):

can i just do m(12m-11)-5 ?

OpenStudy (mathmate):

For your information, the solution to \(12m^2-11m+5 =0\) is \(\large x=\frac{11\pm\sqrt{119}i}{24}\) Perhaps you could discuss this with your prof.

OpenStudy (anonymous):

the answer was : (3m+1)(4m-5) . howw??!1

OpenStudy (mathmate):

When expanded, it gives: 12m^2-11m-5 Check you question. On the other hand, 12m^2-11m-5 =12m^2+4m -15m-5 you can get the grouping by checking product=12*-5=-60 sum=-11 You look for m,n such that m+n=-11, m*n=-60. So, after some trials, you will find m=-15, n=4 Now you can group, using n=4, to get 12m^2+4m -15m-5 Alternatively, using m=-15, you will get 12m^2-15m +4m-5 Both will work in factorization. Note: ratios 12:4=-15:-5, or 12:-15=4:-5